OpenStudy (anonymous):

Find an exact value. sine of {19pi / 12 }

OpenStudy (anonymous):

pi/12=15 degrees 19 .15=285 sin285=? 285=270+15 sin(270+15)=sin270cos15+cos270sin15 = - cos15 = - cos(45-30) = - ( cos45cos30+sin45sin30) = - sqrt(2)/2. sqrt(3)/2 - sqrt(2)/2 .1/2 = - sqrt(6)/4- sqrt(2)/4

OpenStudy (anonymous):

can you fan me

OpenStudy (anonymous):

wow, i actually got it right the first time, thank you!!

OpenStudy (anonymous):

npyw